Instructions for ACL-2005 Proceedings

Hwee Tou Ng Department of Computer Science National University of Singapore

Singapore, 117543

nght@comp.nus.edu.sg

Kemal Oflazer

Faculty of Engineering and Natural Sciences

Sabanci University Istanbul, 34956, Turkey

  oflazer@sabanciuniv.edu

Abstract

This document contains the instructions for preparing a camera-ready manuscript for the proceedings of ACL-2005. The document itself conforms to its own specifications, and is therefore an example of what your manuscript should look like. Authors are asked to conform to all the directions reported in this document.

1 Credits

This document has been adapted from the instructions for ACL-02 proceedings by Eugene Charniak and Dekang Lin, which in turn was based on the formats of earlier ACL and EACL Conference proceedings. Those versions were written by several people, including John Chen, Henry S. Thompson and Donald Walker. Additional elements were taken from the formatting instructions of the International Joint Conference on Artificial Intelligence.

2 Introduction

The following instructions are directed to authors of papers accepted for publication in the ACL-2005 proceedings. All authors are required to adhere to these specifications. Authors are required to provide a Portable Document Format (PDF) version of their papers. The proceedings will be printed on US-Letter paper. Authors from countries in which access to word-processing systems is limited should contact the publication chairs Jason Eisner (jason@cs.jhu.edu) and Philipp Koehn (koehn@csail.mit.edu) as soon as possible.

3 General Instructions

Manuscripts must be in two-column format. Exceptions to the two-column format include the title, authors’ names and complete addresses, which must be centered at the top of the first page, and any fullwidth figures or tables (see the guidelines in Subsection 3.4). Type single-spaced. Start all pages directly under the top margin. See the guidelines later regarding formatting the first page.

The maximum length of a manuscript is eight (8) pages for the main conference, printed singlesided (see Section 4 for additional information on the maximum number of pages).

3.1 Electronically-available resources ACL provides this description in L^{A}TEX2e (acl2005.tex) and PDF format (acl2005.pdf), along with the L^{A}TEX2e style file used to format it (acl2005.sty) and an ACL bibliography style (acl.bst). These files are all available at

http://www.aclweb.org/acl2005/styles/. A

Microsoft Word template file (acl2005.dot) is also available at the same URL. We strongly recommend the use of these style files, which have been appropriately tailored for the ACL-2005 proceedings.

3.2 Format of Electronic Manuscript For the production of the electronic manuscript you must use Adobe’s Portable Document Format (PDF). This format can be generated from postscript files: on Unix systems, you can use ps2pdf for this purpose; under Microsoft Windows, you can use Adobe’s Distiller, or if you have cygwin installed, you can use dvipdf or ps2pdf. Note that some word processing programs generate PDF which may not include all the necessary fonts (esp. tree diagrams, symbols). When you print or create the PDF file, there is usually an option in your printer setup to include none, all or just non-standard fonts. Please make sure that you select the option of including ALL the fonts. Before sending it, test your PDF by printing it from a computer different from the one where it was created. Moreover, some word processor may generate very large postscript/PDF files, where each page is rendered as an image. Such images may reproduce poorly. In this case, try alternative ways to obtain the postscript and/or PDF. One way on some systems is to install a driver for a postscript printer, send your document to the printer specifying “Output to a file”, then convert the file to PDF.

For reasons of uniformity, Adobe’s Times Roman font should be used. In L^{A}TEX2e this is accomplished by putting

\usepackage{times}
\usepackage{latexsym}

in the preamble.

Additionally, it is of utmost importance to specify the US-Letter format (8.5in × 11in) when formatting the paper. When working with dvips, for instance, one should specify -t letter.

Print-outs of the PDF file on US-Letter paper should be identical to the hardcopy version. If you cannot meet the above requirements about the production of your electronic submission, please contact the publication chairs above as soon as possible.

3.3 Layout

Format manuscripts two columns to a page, in the manner these instructions are formatted. The exact dimensions for a page on US-letter paper are:

• Left and right margins: 1in

• Top margin:1in

• Bottom margin: 1in

• Column width: 3.15in

• Column height: 9in • Gap between columns: 0.2in

Papers should not be submitted on any other paper size. Exceptionally, authors for whom it is impossible to format on US-Letter paper, may format for A4 paper. In this case, they should keep the top and left margins as given above, use the same column width, height and gap, and modify the bottom and right margins as necessary. Note that the text will no longer be centered.

3.4 The First Page

Center the title, author’s name(s) and affiliation(s) across both columns. Do not use footnotes for affiliations. Do not include the paper ID number assigned during the submission process. Use the two-column format only when you begin the abstract.

Title: Place the title centered at the top of the first page, in a 15-point bold font. Long title should be typed on two lines without a blank line intervening. Approximately, put the title at 1in from the top of the page, followed by a blank line, then the author’s names(s), and the affiliation on the following line. Do not use only initials for given names (middle initials are allowed). The affiliation should contain the author’s complete address, and if possible an electronic mail address. Leave about 0.75in between the affiliation and the body of the first page.

Abstract: Type the abstract at the beginning of the first column. The width of the abstract text should be smaller than the width of the columns for the text in the body of the paper by about 0.25in on each side. Center the word Abstract in a 12 point bold font above the body of the abstract. The abstract should be a concise summary of the general thesis and conclusions of the paper. It should be no longer than 200 words.

Text: Begin typing the main body of the text immediately after the abstract, observing the twocolumn format as shown in the present document.

Indent when starting a new paragraph. For reasons of uniformity, use Adobe’s Times Roman fonts, with 11 points for text and subsection headings, 12 points for section headings and 15 points for the title. If Times Roman is unavailable, use Computer Modern Roman (L^{A}TEX2e’s default; see section 3.2 above). Note that the latter is about 10% less dense than Adobe’s Times Roman font.

3.5 Sections

Headings: Type and label section and subsection headings in the style shown on the present document. Use numbered sections (Arabic numerals) in order to facilitate cross references. Number subsections with the section number and the subsection number separated by a dot, in Arabic numerals. Do not number subsubsections.

Citations: Citations within the text appear in parentheses as (Gusfield, 1997) or, if the author’s name appears in the text itself, as Gusfield (1997). Append lowercase letters to the year in cases of ambiguities. Treat double authors as in (Aho and Ullman, 1972), but write as in (Chandra et al., 1981) when more than two authors are involved. Collapse multiple citations as in (Gusfield, 1997; Aho and Ullman, 1972).

References: Gather the full set of references together under the heading References; place the section before any Appendices, unless they contain references. Arrange the references alphabetically by first author, rather than by order of occurrence in the text. Provide as complete a citation as possible, using a consistent format, such as the one for Computational Linguistics or the one in the Publication Manual of the American Psychological Association (American Psychological Association, 1983). Use of full names for authors rather than initials is preferred. A list of abbreviations for common computer science journals can be found in the ACM Computing Reviews (Association for Computing Machinery, 1983).

The L^{A}TEX and BibTEX style files provided roughly fit the American Psychological Association format, allowing regular citations, short citations and multiple citations as described above.

Appendices: Appendices, if any, directly follow the text and the references (but see above). Letter them in sequence and provide an informative title: Appendix A. Title of Appendix.

Acknowledgement sections should go as a last section immediately before the references. Do not number the acknowledgement section.

3.6 Footnotes

Footnotes: Put footnotes at the bottom of the page. They may be numbered or referred to by asterisks or other symbols.^{1} Footnotes should be separated from the text by a line.^{2}

3.7 Graphics

Illustrations: Place figures, tables, and photographs in the paper near where they are first discussed, rather than at the end, if possible. Wide illustrations may run across both columns. Do not use color illustrations as they may reproduce poorly.

Captions: Provide a caption for every illustration; number each one sequentially in the form: “Figure 1. Caption of the Figure.” “Table 1. Caption of the Table.” Type the captions of the figures and tables below the body, using 11 point text.

4 Length of Submission

Eight pages (8) is the maximum length of papers for the ACL-2005 main conference. All illustrations, references, and appendices must be accommodated within these page limits, observing the formatting instructions given in the present document. Papers that do not conform to the specified length and formatting requirements are subject to be rejected without review.

Up to two (2) additional pages may be purchased from ACL at the price of $250 per page; please contact the publication chairs above for more information about this option.

References

Alfred V. Aho and Jeffrey D. Ullman. 1972. The Theory of Parsing, Translation and Compiling, volume 1. Prentice-Hall, Englewood Cliffs, NJ.

American Psychological Association. 1983. Publications Manual. American Psychological Association, Washington, DC.

Association for Computing Machinery. 1983. Computing Reviews, 24(11):503–512.

Ashok K. Chandra, Dexter C. Kozen, and Larry J. Stockmeyer. 1981. Alternation. Journal of the Association for Computing Machinery, 28(1):114–133.

Dan Gusfield. 1997. Algorithms on Strings, Trees and Sequences. Cambridge University Press, Cambridge, UK.

1This is how a footnote should appear.

2Note the line separating the footnotes from the text.